OpenStudy (anonymous):

@swagmaster47 and @THEUSER822 plz come help me

OpenStudy (anonymous):

Emilio’s pay varies directly with the number of sales he makes. He made 22 sales and earned $154. How much will he earn for 35 sales?

OpenStudy (anonymous):

154/22 =7 each sale. 35*7=?

OpenStudy (anonymous):

245

OpenStudy (anonymous):

245 dollars

OpenStudy (anonymous):

154/22 =7 each sale. 35*7=245

OpenStudy (anonymous):

yep thats all

OpenStudy (anonymous):

yeah haha

OpenStudy (anonymous):

ok next won ill give yall all testimonees and the person who answers my last quesion gets medal

OpenStudy (anonymous):

A customer paid $4.50 for 6 photo prints. The cost for prints varies directly with the number of prints made. How much will the customer pay for 15 prints?

OpenStudy (anonymous):

wait guys

OpenStudy (unanimoose):

Well first, we need the cost for an individual print.

OpenStudy (unanimoose):

We find that by dividing the cost by the number of given prints. Cost = 4.5 Prints = 6 The individual cost times 15 will give you your answer.

OpenStudy (anonymous):

@Darknight200 i wanna see you try set it up

OpenStudy (anonymous):

dosent have to be right just try